OpenStudy (firejay5):

I need help with Modulos, I wasn't in class when they taught it, so I need a little help. 1. 98mod29 2. 32mod17 3. 17mod 35 4. -47mod45 5. 42mod100 6. 0mod20 7. 52mod26 8. -2mod10 9. 62mod4

OpenStudy (firejay5):

@mathstudent55 @mathmath333

OpenStudy (firejay5):

@dan815

OpenStudy (mathmath333):

u have to find the reaminder when 98 is divided by 29 example \(\large\tt \begin{align} \color{black}{98 (mod 29)\equiv 11}\end{align}\)

OpenStudy (mathmath333):

32 divided by 17 gives remainder 15 \(\large\tt \begin{align} \color{black}{32 (mod17)\equiv 15}\end{align}\)

OpenStudy (firejay5):

don't you mean subtract or add

myininaya (myininaya):

It does get trickier with negative numbers though. So it is best to recall the divison theorem: For any two integers a and b there exist q and r such that a=bq+r where 0<=r<|b| So this means r has to be positive. a mod b should result in a positive integer so for example 13 mod 5=3 since 13=5(2)+3 -13 mod 5=2 since -13=5(-3)+2

OpenStudy (mathmath333):

if there is a negative number u can simply add the divisor to the dividend untill it gets positive example in case of \(-47~~(mod45)\)- \(\large -47+45=-2\implies -2+45=43\) so \(\large\tt \begin{align} \color{black}{-47 (mod45)\equiv 43}\end{align}\) \(\large\tt \begin{align} \color{black}{-2 ~~(mod ~10)\equiv 8-----(-2+10=8)}\end{align}\)

OpenStudy (firejay5):

what do I put on #5. I think the answer is 45

OpenStudy (firejay5):

I meant 42

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ 42~~(mod~100)\equiv 42}\end{align}\) as \(42<100\) \(a~~(mod~b)\equiv a\) \(if~~a<b\)

OpenStudy (firejay5):

@agent0smith #6 0mod20; I think the answer is 0, but I am not for sure if it's the right answer.

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{6.)\\ 0~~(mod~20)\equiv 0}\end{align}\) as i stated \( 0<20\)